OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos (5π/12) I have no clue how to do this....

OpenStudy (anonymous):

the formula is cos=+or- 1+cos/2 all over a square root

OpenStudy (anonymous):

so it will be…. 1 + cos(5pi/12)/2 all under a square root

OpenStudy (lukecrayonz):

@jim_thompson5910

OpenStudy (lukecrayonz):

@timo86m

OpenStudy (anonymous):

that makes absolutely no sense to me...

OpenStudy (lukecrayonz):

@AccessDenied

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

no you need to solve it on you calculator

OpenStudy (anonymous):

ehhh I have an example problem for sin and the answer isn't an exact value and there's a heck of a lot more steps

jimthompson5910 (jim_thompson5910):

cos (5pi/12) = cos ((1/2)*(5pi/6)) cos (5pi/12) = sqrt( (1+cos(5pi/6))/2 ) cos (5pi/12) = sqrt( (1-sqrt(3)/2)/2 ) cos (5pi/12) = sqrt( 1/2-sqrt(3)/4 ) cos (5pi/12) = sqrt( (2-sqrt(3))/4 ) cos (5pi/12) = sqrt(1/4)*sqrt( 2-sqrt(3) ) cos (5pi/12) = (1/2)*sqrt( 2-sqrt(3) )

OpenStudy (anonymous):

Thank you @jim_thompson5910 (:

jimthompson5910 (jim_thompson5910):

yw